OpenStudy (kevin):

1^-4 + 2^-4 + 3^-4 + 4^-4 + ... ------------------------------ = 1^-4 + 3^-4 + 5^-4 + 7^-4 + ... How to solve this?

OpenStudy (raden):

let the Top part is S, and the bottom part is T, so the question is similar to simplifying of S/T

OpenStudy (raden):

S/T = (1^-4 + 2^-4 + 3^-4 + 4^-4 + ...) / (1^-4 + 3^-4 + 5^-4 + 7^-4 + ... )

OpenStudy (raden):

we can grouping into two part of the value of S, S = (1^-4 + 3^-4 + 5^-4 + 7^-4 + ...)+(2^-4 + 4^-4 + 6^-4 + 8^-4 + ...)

OpenStudy (raden):

simplify that fraction and then it can be S/T = 1 + (2^-4 + 4^-4 + 6^-4 + 8^-4 + ...)/(1^-4 + 3^-4 + 5^-4 + 7^-4 + ...)

OpenStudy (raden):

factorize that 2^-4 + 4^-4 + 6^-4 + 8^-4 + ... = 2^-4(1^-4 + 2^-4 + 3^-4 + 4^-4 + ...) = 1/16 (1^-4 + 2^-4 + 3^-4 + 4^-4 + ...)

OpenStudy (raden):

now we have the equation : S/T = 1 + 1/16 (S + T)

OpenStudy (raden):

opssss i mean, it will be S/T = 1 + 1/16 (S+T)/T

OpenStudy (raden):

multiple by 16 to both side give us : 16S/T = 16 + (S + T)/T 16S/T = 16 + S/T + 1 16 S/T - S/T = 17

OpenStudy (kevin):

so the answer is 17?

OpenStudy (raden):

16 S/T - S/T = 17 (16 - 1) S/T = 17 15 S/T = 17 S/T = 17/15

OpenStudy (kevin):

Tq very much